carmen had $130 more than her sister, Rosa. rosa had twice as much money as their youngest sister, Zoe. If their mother gave carmen $60 and rosa $80, Carmen would have twice as much money as rosa. what was the total amount of money they had in the beginning?
I keep getting a negative number.
zoe =x
rosa=2x
carmen= 2x+130
SO....
2(2x+80)=2x+30+60
4x+160=2x+90
4x-2x+160=2x-2x+90
2x+160=90
2x+160-160=90-160
2x=-70
x=-35
what did I do wrong???

Answer by ankor@dixie-net.com(22740)

carmen had $130 more than her sister, Rosa. rosa had twice as much money as their youngest sister, Zoe. If their mother gave carmen $60 and rosa $80, Carmen would have twice as much money as rosa. what was the total amount of money they had in the beginning?
I keep getting a negative number.
zoe =x
rosa=2x
carmen= 2x+130
:
" mother gave carmen $60 and rosa $80, Carmen would have twice as much money as rosa."
Wouldn't the equation be:
2x + 130 + 60 = 2(2x+80)
2x + 190 = 4x + 160
190 - 160 = 2x - 4x
30 = 2x
x = $15 is Zoes amt
then
2(15) = $30 is Rosa's in the beginning
and
2(15) + 130 = $160 is Carmen's original amt
:
:
See if this checks out in the statement
"mother gave carmen $60 and rosa $80, Carmen would have twice as much money as rosa.
160 + 60 = 2(30+80)
:
:
Looking over your work
2(2x+80)=2x+30+60
This should be
2(2x+80) = 2x + 130 + 60; not 30